OpenStudy (anonymous):

how do I verify the identity: cos x ------- = sec x + tan x 1-sin x

OpenStudy (anonymous):

Multiply both top and bottom of fraction by 1+sinx and simplify

OpenStudy (anonymous):

\[\frac{1}{cosx}+\frac{sinx}{cosx}\]

OpenStudy (anonymous):

\[\frac{1+\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

But how do I get to \[\frac{ 1+sin (x) }{ cos (x) }\]?? Do I multiply by the equation by 1+sin and then cos?

OpenStudy (anonymous):

well i took the common factor, like if you want to add 1/2 + 3/2 = 1+3/2, the same method u may use for the above

OpenStudy (anonymous):

okay thank you!

OpenStudy (anonymous):

you're welcome!

OpenStudy (anonymous):

by the way you need to correct the quesion